Fermilab Shutdown Report
7/31/09 – 8/7/09
LINAC
Shutdown Week 8 – On Schedule
- Preaccelerator
- H- Source is operational putting out 56 mA
- Paint job complete of the I- Source & elevator pit
- Started conditioning the I - system
- Low Energy
- Complete RF power turned on in the LRF1-5 & Buncher
- Worked on the LRF4 modulator
- During troubleshooting a capacitor got shorted to ground. It was replaced.
- Worked on the QPS103 power supply
- LE LLRF code upgraded on all the stations
- Ability to tune on beam event specific parameters
- Continue tuning and improving beam transmission to NTF
- High Energy
- Next Week
- Support the Neutron Therapy Facility patient treatment (Tuesday, Thursday, Friday)
- Finalize the HE QPS activities
- Start turning on the Klystrons
- Address problems as they reveal themselves
Booster
Shutdown Week 8 – On Schedule
- Tunnel Activities
- correction element work
- Testing new corrector power supplies
- Checking magnetic fields
- Terminating cables
- All ion pumps are running
- GMPS work continues 90% complete
- East Gallery LCW valved into system
- Booster 95 LCW stable, cleaning up
- Laser Profile Monitor testing
- BRF turn on in progress
- Next Week:
- Continue corrector power supply testing
- Changing door cores (Thursday)
- Safety system tests (Friday)
- GMPS hipot (Friday)
- GMPS commissioning (weekend)
- Power supply turn-on (weekend)
- Beam running on following weekend possibly earlier
- FESS replacing chilled water line BEG
Antiproton Source
Accumulator 4-8GHz DP Kicker Upgrade: ~85% complete
The A50 stochastic kicker tanks are repaired, under vacuum, and we will start the bake on Monday morning. We also developed a vacuum leak in a bellows between the repaired tanks.
Debuncher Septum Magnet Replacement: ~99.5% complete
The repaired magnet is under vacuum and testing will start today. We will pulse the magnet until we are convinced everything is okay. If we run without incident for several months, similar repairs will be done to our spare magnet.
IB7 (AP2 dipole magnet) Replacement: Complete
Other Work
- RF system work. Maintenance work on various RF systems continues
- LCW leak repairs continue
- Target SEM work continues
- A60 cryostat vacuum O-ring replacement complete. When signoffs are done, cryo will sting the tanks and restore LN2. This will be done next week
- When pumping down AP2 after IB7 replacement, a leak was found in an O-ring on the U/S end of the line (radiation damage). The new spool piece has copper gaskets. Pump down of the line is slow but improving
- As time permits, leak testing of parts of both the Accumulator & Debuncher will be done
Main Injector
Tunnel Work
MI-8 Multiwires
- 830, 839 & 840 installed
- 810, 811, 813 & 814 have a systematic failure with the Ti foil and conductive epoxy. We are reverting back to the G10 multiwires. Delay -> 4 days
- 830 may also be affected but will remain installed.
- Two Texas wires also being replace due to vacuum leaks
MI-40
- Pipe fitters are fabricating and installing cooling systems
- We have four of the six kickers being vacuum tested prior to install. Next two are ready for testing. (6 of 7)
- Lead blanket removal on Aug. 10 th followed by Lambertson leak checking/ion pump replacement
- Installed of 2 BPMs @ 403
- Ceiling plates have been installed for shielding
MI eCloud (MI-52)
- Leak tight
- Bake starts tomorrow
- Cloud detectors have been installed
- Plan to finish by vacuum work by Aug. 11 th
Other Work
- Need to install shielding plates at MI-10. Penetrations have been grouted. Sealing leaks today
- MI-30 collimator LVDT investigation continues
- MI-30 collimator mask installation scheduled to begin Aug. 24 th
- MIPS & MIRF work going well, on schedule
- Installing HLS at MI-10.
- Lead blankets: remove from MI-40 on Aug. 10 th
- Lead blankets: remove from MI-10 on Aug. 13 th
- After lead is removed LCW re-hosing will finish at those locations.
- LCW fill set to begin Aug. 18 th and conditioning on Aug. 21 st.
Recycler
Recycler
- Pot RR129 (i.e. MI-10 injection kicker) (90% complete)
- Move 6 magnets ( complete)
- Vacuum maintenance
- Sublimated TSPs between 401 & 409
- Start pump down of MI-10 region on Tuesday?
- Next: Leak check 100-103 (stochastic cooling kickers) & other miscellaneous jobs (IPs, sublimating TSPs)
- Power supplies maintenance
- CIA crate activities
- HLS installation at MI-10
- Almost complete
- Not operational for at least another week though
Electron Cooling
- Should be able to close up Pelletron by August 19
- Desiccant regeneration complete
- 2” GVM ordered
- Last (possibly) outstanding job: check shaft bearings (& replaceif needed)
- Water/Glycol system cleanup and upgrade ( Complete)
- NMR probes replacement (90% complete)
- One last probe needs to be tuned in
Tevatron

Other Items
- Start replacement bad stands 5 of 8 done expect 1 per day
- F1 Hi pot OK ready for cool down
- A1 ready for Hi Pot
- Outside air leak BPM A24-1 will be fixed today then A-2 ready
- Vertical crystal collimator baking install next week
- E-17 flying wire work today
- Will have day and night shift for vacuum work started
- Still to do Dump resistor at F46
- Finish HLS at B0
- Low beta quad alignment at B0 may only be the Q5s and Q6s
NuMI
Shutdown Week 8
- •Replace Target NT-2
- Unable to re-tracked Target from Horn 1
- Z Drive Shaft Coupling was slipping
- Removed Z Drive from Module by Hand
- Unable to Survey Target
- Remove Blue Block and T-Block Shielding
- Slide Target out of Horn 1 with crane
- Move Target Module to Work Cell
- Disconnect Target from Module
- Move Module back to Target Chase
- Replace R-Blocks to Top of Target Pile
- Remove Morgue Covers
- Place Target in Morgue & Replace Covers on Morgue
- Continue Installation of AH Dehumidification System
- Controls System
- Pumps & Heat Exchanger
- EPB Dipoles PEEK Installation Continues
- PEEK replacement Complete on HV101-3, HV101-4, HV101-5
- HV101-4 & 101-5 Installed HV101-3 next in line
- Checkout and Leak Checking of New Texas Multi Wires for PM114 & PMTGT in Process
- There are issues with Vacuum
- Looking at need to bake out Profile Monitors
- Completed Hadron Monitor Replacement
Cryogenic system
TeV at LN2 Temperatures

Tevatron Operations and Maintenance Status
- Continued support AD/MSD for vacuum repairs: D3 is back at LN2 temperature. A1 and A2 under LOTO for MSD. A3 is ready for cooldown but awaiting A2 repairs. E3 and F1 ready for cooldown after instrumentation repairs. D1 is warming up to room temperature for MSD
- LN2 inventory control and controlling magnet temperatures at LN2; feeder 45 power outage last weekend resulted in some relief plates being lifted though were resealed fine
- Expanders group: All Tevatron wet & dry expander overhauls complete; all planned flywheels replaced. CDF & DZero engines are complete. Enormous backlog of shop repair and overhaul work
- Electrical group: Continuing upgrade of Cutler Hummer 120V power. Continuing maintenance of cryo I/O crate, electrical equipment cabinets, various low voltage electrical/electronic maintenance
- Compressor group: Mycom Compr relief valves replaced (5/8). 4/5 IHI has been overhauled
- Ongoing numerous FRIG/Compressor controls/instrumentation maintenance/repairs
---------------------------------------------------------------------------------------
Pending RT houses and relief maintenance, Tevatron cryo is on schedule with purification starting week of 8/24.
CHL HELIUM and NITROGEN PLANTS
- 95% of all scheduled routine mechanical and electrical maintenance is complete
Non-complete critical items:
- Scheduled recycle TSI 2.5MW compressor overhaul for Aug 10 by outside subcontractor
- Expect delivery of repaired turbo expanders for the CHL coldbox-2 on Aug 17. Following by installation of repaired turbines into coldbox-2, test run, and subsequent full cooldown with coldbox-2. Backup plan is to restart with coldbox-1 that operated before shutdown
- Scheduled repairs of compressor D for Aug 10 by outside subcontractor
- Maintenance cleaning of degraded cooling tower for the nitrogen plant completed
- Replacement of the plate & frame CHW/ICW heat exchanger next week
- Repairs of the Nitrogen plant Marley water tower-subcontractor today
-----------------------------------------------------------------------------------------------------
Pending successful TSI and coldbox-2 repairs, maintenance, CHL cryo is on schedule and planning resumption of 24/7 shiftwork on 8/17.
Controls
Completed This Week:
- Network Isolation Test
- Networked DSO issues
- Slowed Kerberos ticket response
- Slowed Windows Domain login response
- Linac startup support
- General Controls Maintenance & Support
Next Week:
- TLG boot Monday
- Support startup activities
- General Controls Maintenance & Support
MiniBooNE
Everything is going well.
MTA
The magnet has been primed.
FESS
Electrical and High Voltage
- Kautz Road SS – buss duct & breaker PMs Work In Progress (WIP)
- Findings – 345kv breaker spring charging mechanism oil leak repairs required. $50k cost & 2 weeks required to order/replace – should still meet KRSS startup of 8/24.
- MI PMs continues & on schedule maintaining schedule changes due to construction & manpower requirements.
- Re-lamping continues with near completion of Switch Yard & the scheduling of repairs
Mechanical
- F3 & E0 pond pump replacements to be scheduled starting next week.
- BGE chilled water 2” line replacement [CUB into BGE pipe chase] to start with tentative 8/14 completion
HVAC
- 90% completed – will return to MI construction areas & MI tunnel de-humidifiers
ICW
- Flushing completed on schedule – Zebra Mussel treatment started 8/7am thru 8/8am
CUB
- 55LCW chiller [#3] tube/header replacement to start 8/10 on should be on line by 8/31
- Chiller cleaning completed on three chillers [1, 4 & 5] - #4 & 5 to be ready for operation end of next week
Switchyard 120 – Work has not started
Collider Detector at Fermilab
- The end plugs will be installed today
- They are transferring to their new backup generator
- Their steel is open and will remain open
D-Zero – On Schedule
Mechanical – On Schedule
- MI blanket removal next week
- MI-10 vacuum work
- Six MI kickers have been built; need one more
Instrumentation – On Schedule
Alignment

Alignment
- Survey of the berm for radiation shielding at MI10 after the civil construction is complete
- Survey of the berm for radiation shielding at MI40 after the civil construction is complete
- Align 7 new kicker magnets to be installed in the MI40 area. See MI Personnel for details. 7 Components
- As-found all the Main Injector and Recycler components in cells102-109. The essential components are: Recycler = magnets and cooling tanks no trims. 40 Components
- As-found all the Main Injector and Recycler components after construction back fill in cells 341-404. Align all kickers and new beam pipe. 40 Components
- Stickmic survey of Muon system including CF and EFs after closing detector. The start and finish dates are preliminary and depend on shutdown progress. Shutdown 2009
- Align 2 EPB dipoles in enclosure B (VH94) before reworked EPB dipoles. Align 2 new dipoles
- Align new permanent magnet to be installed on the Linac side. See MTA personnel for details
Safety
Safety
- Booster re-cored tentatively August 13 th
- Accelerator restart sheets are available from ES&H
Interlock
- Single Line Electrical Drawing Verifications
- P-Bar 95% complete
- Tevatron A Sector started last week
- Audio warning system audio level control instillation
- Booster ESS changes for the new correctors are completed
- Booster configuration control lists updated
- Misc. Tevatron Interlock repairs
- Tevatron ground fault diagnostics installation
- Safety System Testing
- Booster chipmunk testing Thursday August 13th
- Booster enclosure testing Friday August 14th
- MI-8 Line will need to be re-cored and tested and
- Booster CDC tested prior to Booster beam startup
Shutdown Safety
- Injuries
- An employee received frostbite on the right arm area below the elbow from resting it on an ice-encased vent while replacing a seal. First aide case
Continue using the Take Five principles before, during, and after work activities to:
|